# GAMA: High-Performance GEMM Acceleration on AMD Versal ML-Optimized AI Engines

### v2

@@ Abstract
General matrix-matrix multiplication (GEMM) is a fundamental operation in machine learning (ML) applications. We present the first comprehensive performance acceleration of GEMM workloads on AMD’s second-generation AIE-ML (AIE2) architecture, which is specifically optimized for ML applications. Compared to AI-Engine (AIE1), AIE offers increased compute throughput and larger on-chip memory capacity. We propose a novel design that maximizes AIE memory utilization, incorporates custom buffer placement within the AIE2 and staggered kernel placement across the AIE2 array, significantly reducing performance bottlenecks such as memory stalls and routing congestion, resulting in improved performance and efficiency compared to the default compiler provided by AMD. We evaluate the performance benefits of our design at three levels: single AIE, pack of AIEs and the complete AIE array. GAMA achieves state-of-the-art performance, delivering up to 165 TOPS (85% of peak) for int8 precision and 83 TBFLOPS (86% of peak) for bfloat16 precision GEMM workloads. Our solution achieves 8.7%, 9%, 39% and 53.6% higher peak throughput efficiency compared to the state-of-the-art AIE1 frameworks AMA, MAXEVA, ARIES and CHARM, respectively.
@@ I Introduction
AMD Versal is a heterogeneous architecture that integrates Adaptive Intelligent Engines (AIE), programmable logic (PL) and processing system (PS) which consists of CPUs. Such architectures can enable high performance for computationally intensive applications such as machine learning (ML), which are mainly dominated by General Matrix Multiplications (GEMM). While prior research has extensively explored the performance of GEMM on the first-generation AI Engine (AIE1 [3]), no existing work has investigated its performance on the second-generation AIE-ML Engine (AIE2 [5]).
Compared to AIE1, the AIE2111In this paper, for ease of reading, we use the word AIE to refer to AIE2. architecture has larger AIE memory, 2x throughput for int8 precision, and supports bfloat16 precision. Design frameworks from prior work such as MaxEVA [14], AMA [8], CHARM [20] exploit the architecture of AIE1 achieving significant peak performance. These frameworks, however, have architectural dependencies that limit their performance gains to AIE1 and make them not directly adaptable to AIE2.
To efficiently exploit the available enhanced compute throughput of the AIE2, it is crucial to first maximize the single AIE performance and then ensure scaling to maximize the AIE array utilization. At the single AIE level, maximizing performance requires running largest possible GEMM sizes, which in turn requires maximizing AIE memory utilization to fit those sizes. Naively maximizing AIE memory utilization often results in buffer allocation to the nearby AI engines, making it challenging to scale the design to the whole array. Restricting the buffers to a single AIE to facilitate scaling often results in stalls due to the compiler’s default buffer placement strategies. No prior work focuses on buffer placement since none of them tries to maximize AIE memory utilization.
@@
At the single AIE level, we propose a custom buffer placement algorithm that maximizes AIE memory utilization (100% in some cases) and minimizes memory stalls while keeping the buffers in a single AIE. At the pack level, we determine the optimal connectivity and pack size that guarantees scalability and performance, while taking into account the constraints posed by the limited PL-AIE interface resources. Using custom buffer placement to keep buffers within the pack, we make our pack easily scalable. Finally, at the array level we replicate the pack by proposing a staggered placement pattern overcoming the compiler’s routing congestion issues resulting in 94% AIE array utilization. Furthermore, our custom buffer and kernel placement techniques, alleviate the efforts required by the compiler, resulting in a significant reduction in compilation time.
@@
  • GAMA is the first comprehensive study and implementation of GEMM acceleration on the AMD Versal AIE-ML (AIE2) architecture.
@@
  • We open-source GAMA for other researchers to leverage the framework at: (link blinded for review process).
@@ II Related Work
CHARM [20] AIE1 fp32-fp32,
int16-int16,
@@
RSN-XNN [15] AIE1 fp32-fp32 4 384/400 (96%) Yes Manual (Horizontal) - Low (75%)
ARIES [22] AIE1 fp32-fp32,
@@
Since the introduction of AI engines by AMD [2], a wide variety of workloads have been deployed on them [9, 19, 11, 13, 7, 17, 23, 18]. Additionally, several frameworks have mapped and analyzed the performance of GEMM on these AI engines. Table I summarizes the differences between these and GAMA. CHARM [20, 21, 24] accelerated GEMM on AIE1 using a resource conservative approach which enables efficient scaling but results in limited performance. CHARM shows performance for fp32 input and fp32 output (fp32-fp32) [20], int8-int8 [24] and int16-int16 [24] precisions. GAMA maximizes resource utilization (such as AIEs and AIE-PL interfaces) to achieve maximum performance. MaxEVA [14] (fp32-fp32, int8-int32), on the other hand, maximizes GEMM performance while using significant resources on the chip. It uses two different kernels for matrix multiplication and reduction (addition). This limits the overall efficiency of MaxEVA to 80% since only 80% of the AIEs in the design can perform the matrix multiplication operation (20% are used for addition). On the other hand, GAMA uses a single kernel that performs both matrix multiplication and reduction, similar to CHARM. GAMA utilizes 94% of the AIE array for matrix multiplication. It uses a cascade interface for partial sum transfer between AIEs, unlike MaxEVA which uses buffer sharing between nearby engines for reduction. AMA [8] uses a lower precision output (fp32-fp16, int8-int16) improving performance at the cost of potential accuracy loss. MaxEVA and AMA provides AIE-only performance considering that input data is available in PL, unlike CHARM [20], which provides a full implementation including PL and PS. RSN-XNN [15] designs an overlay by proposing a reconfigurable stream network, leveraging the flexibility of FPGA.
Some prior work has focused on the compilation for AIEs. Vyasa [6] extends the Halide [12] DSL compiler to automatically generate code for AIEs, improving programmability. ARIES [22] designs a code generator for AIE1 on Versal and AIE2 on AMD Ryzen processors by leveraging the MLIR framework [10]. Wierse [16] perform performance analysis of the AIE1’s communication interfaces.
Compared to prior work, GAMA is the first framework to show GEMM acceleration on AIE2. GAMA has the highest AIE2 memory utilization with custom buffer placement, a flexible pack size and a staggered kernel placement for efficient scaling.
@@ III Background
The AMD Versal AI Engine is a heterogeneous architecture that combines AI Engines, Programmable Logic (PL), and Processing System (PS). The AI Engines are VLIW vector processors capable of executing multiple instructions per clock cycle. Connected to the AI Engines is an FPGA fabric (or PL) consisting of Lookup Tables (LUTs), flip-flops (FFs), Digital Signal Processing slices (DSPs), Block RAMs (BRAMs), and Ultra RAMs (URAMs). The AI Engines communicate with the PL and DRAM through two interfaces: the PL interface input/output (PLIO) for data transfer between the PL and AI Engines, and NoC interface tiles for data transfer between the AI Engines and DRAM. The AI Engine array architecture provides us with Buffer (an AIE can directly access a neighboring AIE’s memory), Cascade (direct AIE to AIE connectivity) and Via-switch connections to communicate from one AIE to another. The application processor can run a Linux operating system and can be used to control both PL and AI Engines. In this work, we use the 2nd generation of the AI Engine called AIE-ML (or AIE2). The next paragraph discusses the differences compared to AIE1.
@@
Figure 1 illustrates the Versal AIE-ML architecture. At the top are the AIE-ML cores that offer an int8 throughput of 256 MACs/cycle, which is double that of AIE1. Additionally, AIE2 features twice the memory of 64 KB compared to 32 KB in AIE1. Although AIE2 does not support fp32 computations, it incorporates support for bfloat16, contrasting AIE1, which lacks bfloat16 but has fp32 support. AIE2 also includes a wider cascade bus of 512 bits compared to 384 bits in AIE1. Although cascade connectivity remains unidirectional, AIE2 now supports vertical cascading as well. AIE2 also provides us with addition memory storage in the form of memory tiles.
@@ IV GAMA Framework
In this section, we discuss the GAMA framework in detail. We assume that the input/output data is available in the PL and use the PLIO interface for communication. We use AMD’s Versal VE2802 device which has 304 engines in a grid of 8 rows x 38 columns. This device has 112 input PLIOs (PL to AIE) and 84 output PLIOs (AIE to PL). Each PLIO is configured to be 128 bit wide.
@@ IV-A Single AIE kernel design
Kernel size: Our kernels are designed using AMD APIs leveraging the generic (templated) matrix multiplication kernel from AMD’s user API guide [1], thereby ensuring their portability across a variety of Versal devices. We support multiple sizes, i.e., the M, K, and N matrix dimensions, for a single kernel. Although our kernel code supports variety of M, K and N sizes, selecting the right size is crucial for high performance and AIE kernel compute utilization. The kernel size selection is governed by the input/output precision, the AIE memory limit and the balance between compute and PL-AIE communication, which is defined by the compute-to-communication ratio $\gamma$. Equation 1 calculates the theoretical kernel compute cycles based on the workload dimension and the peak throughput of the chip for a specific precision. Equations 2, 3 and 4 calculate the cycles required to transfer the input matrix A and matrix B from PL to AIE and the output matrix C from AIE to PL, where $sizeof()$ depends on the precision used. Each AIE has two input PLIO and one output PLIO connections. This enables read, compute, and write to be performed at the same time in a pipelined manner. Equation 6 shows the calculation for $\gamma$. A ratio $<1$ indicates that the workload is PLIO bandwidth bound and $>1$ indicates that it is compute bound.
@@
The internal memory of the AIE also presents a constraint on the size of M, K and N as shown by Equation 7. Since double or ping-pong buffering is used to overlap computation in AIE and the AIE-PL communication, the memory requirement is doubled.
@@
We perform a custom buffer address placement using a simple algorithm that maximizes AIE memory utilization while minimizing stalls. Our custom buffer address placement is guided by the set of rules that determine the buffer location based on the matrix type (input and output) and buffer category (Ping or Pong). Our rules are as follows: (a) Never assign ping and pong buffers of the same matrix to the same bank. (b) Never assign ping and pong buffers of the same matrix to the adjacent bank. (c) Always assign the buffer of matrix A and B to a different bank. Algorithm 1 encapsulates these rules and automates buffer address placement for various matrix sizes and precisions. It performs a simple exhaustive search and generates buffer addresses that satisfy the rules mentioned above. It takes the dimensions of matrix M, K, N, and the input and output precision as inputs and first calculates the size of buffers A, B, and C (lines 1-4). If the sum of the sizes of the buffer exceeds 64 KB then it exits (Line 5). A buffer list is initialized for all the buffers that need address allocation (Line 7). The order of the list is important. We iterate over all banks of to check if the corresponding buffer is suitable for allocation (Lines 9-31). Banks have two spots for buffers, which means that max two buffers can fit in one bank. Matrix A and B are only allocated when there are two spots available in the bank and the adjacent banks do not contain the corresponding double buffers (Lines 12-13). If the constraint is satisfied, the buffer is placed in the bank, and a start address is allocated (lines 15-18). Matrix C can be placed as the second buffer in the banks that contain matrix A or B (Line 20). The matix C buffer is placed and a start address is assigned to the buffer (Lines 21-25). The sum of both buffers can exceed the capacity (16 KB) of the bank, thus in this scenario the start address of the other buffer in the next bank is shifted by the offset (Lines 27-29). Once finished, the algorithm returns the stating addresses of all six buffers.
@@
3:Initialize: aie_mem.size = 65536 aie_mem.banks = 4
4:buf_A, buf_B, buf_C $\leftarrow$ $M\times K\times ip\_p$, $K\times N\times ip\_p$, $M\times N\times op\_p$
5:if check_overflow(M, K, N, ip_p, op_p, aie_mem.size) then
6:  exit
7:buf_list[:] $\leftarrow$ ping_A,pong_A,ping_B,pong_B,ping_C,pong_C
8:banks $\leftarrow$ aie_mem.banks
9:for buf in buf_list do
10:  for bank in range(banks) do
11:   if  $\text{buf}\in(Mat_{A}orMat_{B})$ then
@@
13:      continue
14:    else
15:      banks[bank].buffer $\leftarrow$ buf
16:      banks[bank].free_space =-=- = buf.size
17:      banks[bank].free_spots --- -
18:      buf.start_addr $\leftarrow$ banks[bank].start_addr
19:   else if  $\text{buf}\in(Mat_{C})$ then
20:    if banks[bank].free_spot $>$then
21:      banks[bank].buffer $\leftarrow$ buf
22:      banks[bank].free_space =-=- = buf.size
23:      if banks[bank].free_spots $=$then
24:       buf.start_addr $\leftarrow$ banks[bank].start_addr
25:      if banks[bank].free_spots $=$then
26:       buf.start_addr $\leftarrow$ banks[bank].start_addr + banks[bank].buffer[0].size
27:      if banks[bank].free_space $<$then
28:       offset $\leftarrow$ (banks[bank].buffer[0].size + banks[bank].buffer[1].size) - banks[bank].size
29:       banks[bank+1].buffer.start_addr $\leftarrow$ banks[bank+1].start_addr + offset
30:       banks[bank].free_spots--- -
31:      break
32:return buf_list.addr

### v3

@@ Abstract
General matrix-matrix multiplication (GEMM) is a fundamental operation in machine learning (ML) applications. We present the first comprehensive performance acceleration of GEMM workloads on AMD’s second-generation AIE-ML architecture, which is specifically optimized for ML applications. Compared to AI-Engine (AIE), AIE-ML offers increased compute throughput and larger on-chip memory capacity. We propose a novel design that maximizes AIE-ML memory utilization, incorporates custom buffer placement within the AIE-ML and staggered kernel placement across the AIE-ML array, significantly reducing performance bottlenecks such as memory stalls and routing congestion, resulting in improved performance and efficiency compared to the default AMD’s compiler. We evaluate the performance benefits of our design at three levels: single AIE-ML, pack of AIE-ML’s and the complete AIE-ML array. GAMA achieves state-of-the-art performance, delivering up to 165 TOPS (85% of peak) for int8 precision and 83 TBFLOPS (86% of peak) for bfloat16 precision GEMM workloads. Our solution achieves 8.7%, 9%, 39% and 53.6% higher peak throughput efficiency compared to the state-of-the-art AIE frameworks AMA, MAXEVA, ARIES and CHARM, respectively.
@@ I Introduction
AMD Versal is a heterogeneous architecture that integrates Adaptive Intelligent Engines (AIE), programmable logic (PL) and processing system (PS) which consists of CPUs. Such architectures can enable high performance for computationally intensive applications such as machine learning (ML), which are mainly dominated by General Matrix Multiplications (GEMM). While prior research has extensively explored the performance of GEMM on the first-generation AI Engine (AIE) [3], no existing work has investigated its performance on the second-generation AIE-ML Engine [5].
Compared to AIE, the AIE-ML [^1] In this paper, for ease of reading, we use the word AIE to refer to AIE-ML. architecture has larger AIE memory, 2x throughput for int8 precision, and supports bfloat16 precision. Design frameworks from prior work such as MaxEVA [14], AMA [8], CHARM [20] exploit the architecture of AIE achieving significant peak performance. These frameworks, however, have architectural dependencies that limit their performance gains to AIE and make them not directly adaptable to AIE-ML.
To efficiently exploit the available enhanced compute throughput of the AIE-ML, it is crucial to first maximize the single AIE performance and then ensure scaling to maximize the AIE array utilization. At the single AIE level, maximizing performance requires running largest possible GEMM sizes, which in turn requires maximizing AIE memory utilization to fit those sizes. Naively maximizing AIE memory utilization often results in buffer allocation to the nearby AI engines, making it challenging to scale the design to the whole array. Restricting the buffers to a single AIE to facilitate scaling often results in stalls due to the compiler’s default buffer placement strategies. No prior work focuses on buffer placement since none of them tries to maximize AIE memory utilization.
@@
At the single AIE level, we propose a custom buffer placement algorithm that maximizes the AIE memory utilization (100% in some cases) and minimizes memory stalls while keeping the buffers in a single AIE. At the pack level, we determine the optimal connectivity and pack size that guarantees scalability and performance, while taking into account the constraints posed by the limited PL-AIE interface resources. Using custom buffer placement to keep buffers within the pack, we make our pack easily scalable. Finally, at the array level we replicate the pack by proposing a staggered placement pattern overcoming the compiler’s routing congestion issues resulting in 94% AIE array utilization. Furthermore, our custom buffer and kernel placement techniques, alleviate the efforts required by the compiler, resulting in a significant reduction in compilation time.
@@
  • GAMA is the first comprehensive study and implementation of GEMM acceleration on the AMD Versal AIE-ML architecture. It adopts a scalability-first approach that emphasizes scalability over the performance of individual kernels, promoting regularity in kernel mapping to facilitate straightforward replication.
@@
  • We open-source GAMA for other researchers to leverage the framework at: https://github.com/advent-lab/GAMA.
@@ II Related Work
CHARM [20] AIE fp32-fp32 4 (fp32) 384/400 (96%) Yes Manual (Horizontal) Custom Low (75%)
AutoMM [24] AIE int16-int16,
@@
RSN-XNN [15] AIE fp32-fp32 4 384/400 (96%) Yes Manual (Horizontal) - Low (75%)
ARIES [22] AIE fp32-fp32,
@@
Since the introduction of AI engines by AMD [2], a wide variety of workloads have been deployed on them [9, 19, 11, 13, 7, 17, 23, 18]. Additionally, several frameworks have mapped and analyzed the performance of GEMM on these AI engines. Table I summarizes the differences between these and GAMA. CHARM [20, 21] and AutoMM [24] accelerated GEMM on AIE using a resource conservative approach which enables efficient scaling but results in limited performance. CHARM and AutoMM show performance for the fp32 input and fp32 output (fp32-fp32) [20] and int8-int8 and int16-int16 [24] precisions, respectively. GAMA maximizes resource utilization (such as AIEs and AIE-PL interfaces) to achieve maximum performance. MaxEVA [14] (fp32-fp32, int8-int32), on the other hand, maximizes GEMM performance while using significant resources on the chip. It uses two different kernels for matrix multiplication and reduction (addition). This limits the overall efficiency of MaxEVA to 80% since only 80% of the AIEs in the design can perform the matrix multiplication operation (20% are used for addition). On the other hand, GAMA uses a single kernel that performs both matrix multiplication and reduction, similar to CHARM. GAMA utilizes 94% of the AIE array for matrix multiplication. It uses a cascade interface for partial sum transfer between AIEs, unlike MaxEVA which uses buffer sharing between nearby engines for reduction. AMA [8] uses a lower precision output (fp32-fp16, int8-int16) improving performance at the cost of potential accuracy loss. MaxEVA and AMA provides AIE-only performance considering that input data is available in PL, unlike CHARM [20], which provides a full implementation including PL and PS. RSN-XNN [15] designs an overlay by proposing a reconfigurable stream network, leveraging the flexibility of FPGA.
Some prior work has focused on the compilation for AIEs. Vyasa [6] extends the Halide [12] DSL compiler to automatically generate code for AIEs, improving the programmability. ARIES [22] designs a code generator for AIE on Versal and AIE-ML on AMD Ryzen processors by leveraging the MLIR framework [10]. The ARIES framework supports AIE-ML on Ryzen CPUs, demonstrating performance results for ResNet instead of GEMM. Therefore, a direct comparison with ARIES is not feasible unless modifications are made to enable support for Versal devices. Wierse [16] perform performance analysis of the AIE’s communication interfaces.
Compared to prior work, GAMA is the first framework to show GEMM acceleration on AIE-ML. GAMA has the highest AIE memory utilization with custom buffer placement, a flexible pack size and a staggered kernel placement for efficient scaling.
@@ III Background
The AMD Versal AI Engine is a heterogeneous architecture that combines AI Engines, Programmable Logic (PL), and Processing System (PS). The AI Engines are VLIW vector processors capable of executing multiple instructions per clock cycle. Connected to the AI Engines is an FPGA fabric (or PL) consisting of Lookup Tables (LUTs), flip-flops (FFs), Digital Signal Processing slices (DSPs), Block RAMs (BRAMs), and Ultra RAMs (URAMs). The AI Engines communicate with the PL and DRAM through two interfaces: the PL interface input/output (PLIO) for data transfer between the PL and AI Engines, and NoC interface tiles for data transfer between the AI Engines and DRAM. The AI Engine array architecture provides us with Buffer (an AIE can directly access a neighboring AIE’s memory), Cascade (direct AIE to AIE connectivity) and Via-switch connections to communicate from one AIE to another. The application processor can run a Linux operating system and can be used to control both PL and AI Engines. In this work, we use the 2nd generation of the AI Engine called AIE-ML. We further discuss the differences compared to AIE.
@@
Figure 1 illustrates the Versal AIE-ML architecture. At the top are the AIE-ML cores that offer an int8 throughput of 256 MACs/cycle, which is double that of AIE. Additionally, AIE-ML features twice the memory of 64 KB compared to 32 KB in AIE. Although AIE-ML does not support fp32 computations, it incorporates support for bfloat16, contrasting AIE, which lacks bfloat16 but has fp32 support. AIE-ML also includes a wider cascade bus of 512 bits compared to 384 bits in AIE. Although cascade connectivity remains unidirectional, AIE-ML now supports vertical cascading as well. AIE-ML also provides us with addition memory storage in the form of memory tiles.
@@ IV GAMA Framework
In this section, we discuss the GAMA framework in detail. We assume that the input/output data is available in the PL and use the PLIO interface for communication. We use AMD’s Versal VE2802 device which has 304 engines in a grid of 8 rows x 38 columns. This device has 112 input PLIOs (PL to AIE) and 84 output PLIOs (AIE to PL). Each PLIO is configured to be 128 bit wide. Matrix multiplication workload is specified with the notation M, K, and N. Where, matrix A has dimensions M $\times$ K, matrix B has dimensions K $\times$ N, and the resulting matrix has dimensions M $\times$ N.
@@ IV-A Single AIE kernel design
Kernel size: Our kernels are designed using AMD APIs leveraging the generic (templated) matrix multiplication kernel from AMD’s user API guide [1], thereby ensuring their portability across a variety of Versal devices. We support multiple sizes, i.e., the M, K, and N matrix dimensions, for a single kernel. Although our kernel code supports variety of M, K and N sizes, selecting the right size is crucial for high performance and AIE kernel compute utilization. The kernel size selection is governed by the input/output precision, the AIE memory limit and balance between compute and PL-AIE communication, which is defined by the compute-to-communication ratio $\gamma$. Equation 1 calculates the theoretical kernel compute cycles based on the workload dimension and the peak throughput of the chip for a specific precision. Equations 2, 3 and 4 calculate the cycles required to transfer the input matrix A and matrix B from PL to AIE and the output matrix C from AIE to PL, where $sizeof()$ depends on the precision used. Each AIE has two input PLIO and one output PLIO connections. This allows simultaneous pipelined read, compute, and write operations. Equation 6 shows the calculation for $\gamma$. A ratio <1<1 means the workload is PLIO bandwidth bound, while >1>1 means it is compute bound.
@@
The internal memory of the AIE presents a constraint on the size of M, K and N as shown by Equation 7. Since ping-pong buffering is used to overlap computation in AIE and the AIE-PL communication, the memory requirement is doubled.
@@
We perform a custom buffer address placement using a simple algorithm that maximizes AIE memory utilization while minimizing stalls. Our custom buffer address placement is guided by the set of rules that determine the buffer location based on the matrix type (input and output) and buffer category (Ping or Pong). Our rules are as follows: (a) Never assign ping and pong buffers of the same matrix to the same bank. (b) Never assign ping and pong buffers of the same matrix to the adjacent bank. (c) Always assign the buffer of matrix A and B to a different bank. Algorithm 1 encapsulates these rules and automates buffer address placement for various matrix sizes and precisions. It performs a simple exhaustive search and generates buffer addresses that satisfy the rules mentioned above. It takes the dimensions of matrix M, K, N, and the input and output precision as inputs and first calculates the size of buffers A, B, and C (lines 1-5). If the sum of the sizes of the buffer exceeds 64 KB then it exits (Line 6). A buffer list is initialized for all the buffers that need address allocation (Line 8). The order of the list is important. We iterate over all banks of to check if the corresponding buffer is suitable for allocation (Lines 10-32). Banks have two spots for buffers, which means that max two buffers can fit in one bank. Matrix A and B are only allocated when there are two spots available in the bank and the adjacent banks do not contain the corresponding double buffers (Lines 13-14). If the constraint is satisfied, the buffer is placed in the bank, and a start address is allocated (lines 16-19). Matrix C can be placed as the second buffer in the banks that contain matrix A or B (Line 21). The matix C buffer is placed and a start address is assigned to the buffer (Lines 22-26). The sum of both buffers can exceed the capacity (16 KB) of the bank, thus in this scenario the start address of the other buffer in the next bank is shifted by the offset (Lines 28-30). Once finished, the algorithm returns the stating addresses of all six buffers.
@@
3:Initialize: aie_mem.size = 65536 aie_mem.bnks = 4
4:buf_A, buf_B $\leftarrow$ $M\times K\times ip\_p$, $K\times N\times ip\_p$
5:buf_C $\leftarrow$ $M\times N\times op\_p$
6:if chk_ofl(M, K, N, ip_p, op_p, aie_mem.size) then
7:  exit
8:buf_list[:] $\leftarrow$ ping_A,pong_A,ping_B,pong_B,ping_C,pong_C
9:bnks $\leftarrow$ aie_mem.bnks
10:for buf in buf_list do
11:  for b in range(bnks) do
12:  if  $\text{buf}\in(Mat_{A}orMat_{B})$ then
@@
14:   continue
15:    else
16:   bnks[b].buffer $\leftarrow$ buf
17:   bnks[b].free_space =-= buf.size
18:   bnks[b].free_spots --
19:   buf.start_addr $\leftarrow$ bnks[b].start_addr
20:  else if  $\text{buf}\in(Mat_{C})$ then
21:    if bnks[b].free_spot >>then
22:   bnks[b].buffer $\leftarrow$ buf
23:   bnks[b].free_space =-= buf.size
24:   if bnks[b].free_spots ==then
25:    buf.s_addr $\leftarrow$ bnks[b].s_addr
26:   if bnks[b].free_spots ==then
27:    buf.s_addr $\leftarrow$ bnks[b].s_addr + bnks[b].buffer[0].size
28:   if bnks[b].free_space <<then
29:    offset $\leftarrow$ (bnks[b].buffer[0].size + bnks[b].buffer[1].size) - bnks[b].size
30:    bnks[b+1].buffer.s_addr $\leftarrow$ bnks[b+1].s_addr + offset
31:    bnks[b].free_spots--
32:   break
33:return buf_list.addr